OpenStudy (anonymous):

Determine the equation of the line that passes through the points (2, -5) and (7, 0). Use the form y = mx + b.

OpenStudy (anonymous):

Find the slope first, using \[\huge m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] Once you have that use the point - slope form. \[\huge y-y _{1}=m(x-x _{1})\] Form that into y = mx+b

OpenStudy (anonymous):

m = slope (x1,y1) your first point.

OpenStudy (anonymous):

y = 2x - 5 ?

OpenStudy (anonymous):

y = x + 7 ???

OpenStudy (anonymous):

You sure it's +7?

OpenStudy (anonymous):

A. y = 2x - 5 B. y = x + 7 C. y = 5x + 5 D. y = x - 7

OpenStudy (anonymous):

y=x-7

OpenStudy (anonymous):

Correct ?

OpenStudy (anonymous):